The force P is sufficient to accelerate the three blocks as a unit along a frictionless surface, where mC>mB> mA. If FAB is the force that block A exerts on block B and FBC is the force block B exerts on block C, rank the magnitudes of the three forces from greates to least.The force P is sufficient to accelerate the three
Answer #1

Here,

let the acceleration of the blocks is a

Now, for FAB

Using second law of motion

FAB = (mC + mB) * a

for FBC

FBC = mC * a

hence , for force P = (mA + mB + mC) * a

P > FAB > FBC